Lucifer February Com Night 5 February 10-11 2009

Observers: N.Ageorges, W.Seifert, V.Knierem (LUCIFER), DThompson
Telescope Operator: W.Wack
SW Support:
Telescope Support: DOfficer, JMorris


Ice still on dome, enough to prevent opening even though it is clear outside. +10 degree change in outside temp, -96% change in humidity.

Partial night, TRACK mode testing by LUCIFER with the dome closed.

UT 11:01:25.5 Offset sent from LUCIFER returned error. It looks like the same problem that occurred yesterday, in IT#2002. However, tonight we also have the rotator claiming bad polynomials. The suggested fix of switching the mount to velocity mode and then back to track did not clear the problem. We stopped and restarted PCS (~04:13 MST) to clear the scrolling bad poly messages and resent the preset, which succeeded.

-- DavidThompson - 11 Feb 2009
Topic revision: r2 - 11 Feb 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback